# UniVS: Unified and Universal Video Segmentation with Prompts as Queries

### v1

@@ 1 Introduction
Video segmentation (VS) partitions a video sequence into different regions or segments, facilitating many applications such as video understanding, region-guided video generation, interactive video editing and augmented reality, etc. VS tasks can be divided into two groups: category-specified VS and prompt-specified VS. The former focuses on segmenting and tracking entities from a predefined set of categories. Typical tasks along this line include video instance [yang2019video, qi2021occluded], semantic [vspw] and panoptic segmentations [vps, vipseg] (VIS/VSS/VPS), where the object category information needs to be specified. Another group focuses on identifying and segmenting specific targets throughout the video, where visual prompts or textual descriptions of the targets need to be provided. Visual prompts can be points, bounding boxes, masks, or scribbles. Prompt-specified VS tasks include video object segmentation (VOS) [Perazzi2016vos, youtubevos, mose, burst], Panoptic VOS (PVOS) [xu2023viposeg] and referring VOS (RefVOS) [refvos]. Each of above VS tasks has established its own protocol for dataset annotation, as well as model evaluation.
VS has a close relationship with image segmentation (IS) [everingham2010pascalvoc, lin2014microsoft, ade20k, panoptic_seg, cordts2016cityscapes, xu2016deep_interactive_seg, extreme_interactive_seg, refcoco, zhou2023interseg-GP], which has similar types of segmentation tasks to VS. In the past decades, the model performance on each individual IS/VS task has been significantly improved, and many well-known network architectures have been developed [he2017mask, bolya2019yolact, chen2017deeplab, fcn_ss, 2015Unet, deeperlab_ps, panoptic_deeplab, extreme_interactive_seg, refer, grounding_refer, Li_2021_CVPR, wang2020vistr, xmem, IDOL, huang2022minvis, openvis]. Some architectures [k-net, cheng2021mask2former, li2022maskdino, zou2022xdecoder, videoknet, videokmax, tubelink, dvis] have also been proposed to deal with multiple segmentation tasks; however, these architectures require separate training and inference for different tasks because of the different annotations and evaluation protocols. Fortunately, the recent advancements in vision-language models [li2022blip, clip, 12-in-1, li2021glip, regionclip, liu2023llava, lin2023videollava, liu2023LargeWorldModel, yu2023magvit] align and harmonize the multimodal feature representations, which bridge the labeling gaps across various IS/VS tasks. As a result, some unified segmentation models [freeseg, seem, semanticsam] have emerged to process multiple segmentation tasks simultaneously, which can be jointly trained on different datasets and tasks. These methods generally convert prompt-guided target segmentation into the category-specified segmentation problem [he2017mask, zhu2020deformable_detr, cheng2021mask2former], which first predicts masks for all potential entities per frame and then utilizes a post-processing matching to find the target. For example, SEEM [seem] captures prompt information of the target via concatenating learnable queries and prompt features as keys and values in the self-attention layer of the mask decoder, showing good versatility in various IS tasks.
@@
Compared to IS, the additional challenge of VS lies in the temporal consistency [wang2021unitracker] of segmentation across frames in a video sequence. Existing unified models [uninext, tarvis] for VS tasks are mostly inspired by the unified IS models. They segment the video sequence frame by frame, and then use a similarity matching step to associate common objects or find the targets for category-specified and prompt-specified VS tasks, respectively. For example, UNINEXT [uninext] is specially designed for object-centric IS and VS tasks. Though UNINEXT performs very well in several aspects, it becomes ineffective when segmenting ‘stuff’ entities, such as ‘sky’. To adapt to different VS tasks, TarVIS [tarvis] subdivides the learnable queries within the mask decoder into four groups: semantic queries, class-agnostic instance queries, object queries and background queries. However, TarVIS falls short in encoding linguistic information, hindering it from resolving language-guided VS tasks.
@@
To alleviate the above issues, we propose a novel unified VS architecture, namely UniVS, by using prompts as queries. For each target of interest, UniVS averages the prompt features from previous frames as its initial query. A target-wise prompt cross-attention (ProCA) layer is introduced in the mask decoder to integrate comprehensive prompt features stored in the memory pool. The initial query and the ProCA layer play a crucial role in the explicit and accurate decoding of masks. On the other hand, by taking the predicted masks of entities from previous frames as their visual prompts, UniVS can convert different VS tasks into the task of prompt-guided target segmentation task, eliminating the heuristic inter-frame matching. The overall process of UniVS is depicted in Fig. 2(b). UniVS not only unifies the different VS tasks (see Fig. UniVS: Unified and Universal Video Segmentation with Prompts as Queries) but also naturally achieves universal training and testing, resulting in robust performance across different scenarios. It shows a commendable balance between performance and universality on 10 challenging VS benchmarks, covering VIS, VSS, VPS, VOS, RefVOS and PVOS tasks. To the best of our knowledge, UniVS is the first work which can unify all the existing VS tasks successfully in a single model.
@@ 2 Related Work
Category-specified VS tasks includes video instance [yang2019video], semantic [vspw] and panoptic segmentation [vipseg] tasks, which aim to correctly partition the regions and label them with specific categories or open-world vocabulary[openseed, li2023opensd, openvis]. Building on representative query-based image detection and segmentation methods [carion2020detr, zhu2020deformable_detr, lishuai2023one, cheng2021mask2former], recent approaches [Li_2021_CVPR, wang2020vistr, seqformer, IDOL, heo2022vita, huang2022minvis, li2023mdqe, GenVIS, tubelink, tubeformer, videoknet] focus on efficiently encoding both short-term and medium-term temporal information. Some methods [seqformer, IDOL, huang2022minvis, li2023mdqe, panoptic_fcn] utilize the consistency of the relative relationship between objects in a short period of time to associate entities across frames, while the latest state-of-the-arts [heo2022vita, GenVIS, videoknet, tubelink, dvis] design learnable trackers between multiple video clips to better learn entity motion information over short to medium time periods, thus maintaining more temporally consistent segmentation results. This continual learning strategy also bridges the gap between training and inference.
Prompt-specified VS tasks consist of video object segmentation (VOS) [youtubevos, davis, mose, burst], panoptic VOS [xu2023viposeg] and referring VOS [refvos], which aim to re-identify and segment the target objects with visual/text prompts in a video. Recent offline models [xmem, yang2022deaot, referformer, grounding_refer, li2022r2vos, Miao_2023_ICCV_sgmg, cheng2023cutie] focus more on designing long-term information propagation modules to transfer previous image features and corresponding masks to the target frame to predict masks. This helps more precisely identify and track the movement trajectory of objects throughout the entire video sequence (more than 200 frames), but also makes the network architecture heavy [yang2022deaot, Wu_2023_ICCV_uniref]. Due to the slow speed of offline models, online models [wu2023onlinerefer, jointformer_vos, Wu_2023_ICCV_uniref] come back into focus, aiming to keep the right balance between speed and performance.
Unified VS models [cheng2021mask2former-video, videoknet, videokmax, tubelink, dvis, uninext, tarvis, cheng2023DEVA_tracking, Wu_2023_ICCV_uniref] have emerged recently. TubeLink [tubelink] and DVIS [dvis] employ a single framework to accomplish different category-specified VS tasks, but they still need to be trained separately for each task and cannot be used for prompt-specified VS tasks. UniRef [Wu_2023_ICCV_uniref] uses a model to unify prompt-specified VS tasks, but it cannot handle category-specified VS tasks. To handle more types of VS tasks, TarVIS [tarvis] decouples learnable queries of mask decoder into four groups: semantic, class-agnostic instance, object and background queries. However, TarVIS fails to solve RefVOS due to the lack of language encoding capability. UNINEXT [uninext] is a unified object-centric segmentation model on images and videos, demonstrating universal applicability in several aspects. Unfortunately, UNINEXT cannot handle entities with ‘stuff’ categories.
@@ 3 Methodology
UniVS contains of three main modules: the Image Encoder, the Prompt Encoder and the Unified Video Mask Decoder, as depicted in Fig. 3. The Image Encoder transforms the input RGB images to the feature tokens, while the Prompt Encoder translates raw visual/text prompts into the prompt embeddings. The Unified Video Mask Decoder explicitly decodes masks for any entity or prompt-guided target in the input video.
@@ 3.1 Image and Prompt Encoders
Image encoder contains a backbone and a pixel decoder [cheng2021mask2former]. The backbone maps the RGB image $X\in R^{3\times H\times W}$ into multi-scale features, and the pixel decoder further fuses features across scales to enhance the representation: $\{F_{s}\}_{s=1}^{4}=\text{ImEnc}(X),$ where $F_{s}\!\in\!R^{C\times H_{s}\times W_{s}}$ is the $s$-th scale image embeddings, $H_{s},W_{s}$ are its height and width and $C$ is the number of channels. The resolutions of multi-scale feature maps are 1/32, 1/16, 1/8 and 1/4 of that of the input image, respectively.
Prompt encoder converts the input visual/text prompts into the prompt embeddings. The visual prompts can be clicked points, boxes, masks and scribbles, etc. For the $i$-th target, we use $X_{i}^{*}\in R^{l_{i}\times 3}$ and $y_{i}^{*}\in R^{l_{i}}$ to represent its prompt-specified pixels and corresponding segment IDs, where $l_{i}$ is the total number of prompt-specified pixels. To convert the visual prompts into image embeddings, we adopt the Visual Sampler strategy proposed in SEEM [seem]. It samples $l^{*}$ points from the prompt-specified pixels for each target, and extracts point features from the 3rd scale image embeddings $F_{3}$ as its visual prompt embeddings:
@@
The language prompt can be category names, such as ‘person’, and textual expressions, such as ‘a person is skateboarding’. Following [devlin2018bert, clip], we feed a category name or an expression into a tokenizer to get its string tokens $L_{i}^{*}\in R^{l^{*}\times C^{t}}$, which are input into the CLIP text encoder to obtain the text embeddings. We then introduce a single cross-attention layer to achieve language-image embeddings interaction, where the query is text embeddings, the keys and values are flattened multi-scale image embeddings $F\in R^{C\times(\sum_{s=1}^{3}H_{s}W_{s})}$. This process can be formulated as:
@@ 3.2 Unified Video Mask Decoder
where $P_{i}^{*}$ and $y_{i}^{*}$ are prompt features and associated segment IDs for the $i$-th target, and $M_{i}^{t}$ is its predicted masks in the $t$-th frame. $N_{p}$ and $V$ are the total numbers of provided targets and frames in the video, respectively. To achieve our goal, we adapt the mask decoder of Mask2Former [cheng2021mask2former, cheng2021mask2former-video], which was initially designed for generic segmentation tasks with a set of learnable queries, by introducing a side stream that takes the mean of prompt features as input queries. As illustrated in the right yellow area of Fig. 3, our proposed unified video mask decoder comprises four key components: target-wise prompt cross-attention layer, an image cross-attention layer, a separated self-attention layer and a feed-forward network (FFN, which is omitted in Fig. 3).
@@ 4 Training and Inference
4.1 Training Losses and Stages
Training Losses. There are three terms in the training loss:
where $\lambda_{\text{mask}},\lambda_{\text{cls}}$ and $\lambda_{\text{reid}}$ are hyper-parameters to balance the loss terms. Their default values are set to 5, 3, 0.5, respectively. The commonly used Dice loss [dice1945dice] and BCE loss [jadon2020semseg_loss_survey] are employed for pixel-wise mask supervision. The classification loss [lin2017focalloss] only applies to category-specified VS tasks, which utilizes the similarity between text embeddings and query embeddings as the score matrix for label assignment. The contrastive loss and the auxiliary loss [pang2021quasi] are used as the ReID loss, which encourages the samples of the same target to be closer while pushing samples of different targets away from each other.
Training Stages. The training process of UniVS consists of three stages: image-level training, video-level training and long video fine-tuning. In the first stage, UniVS is trained on multiple image segmentation datasets, pretraining the model with image-level annotations for a good visual representation. In the second stage, we feed a short clip of three frames to the pretrained model, and fine-tune it on video segmentation datasets to perceive entity changes over a short period of time. In the third stage, we employ long video sequences of more than five frames to further fine-tune the unified video mask decoder, encouraging it to learn more discriminative features and trajectory information over a longer time period. To optimize memory usage, we freeze the backbone weights in the last two stages and further freeze the pixel decoder in the final stage. In each iteration, all samples within a batch come from the same dataset. We found that this sampling strategy can make the training more stable compared with the mixed sampling from different datasets.
@@ 4.2 Unified Streaming Inference Process
For prompt-specified VS tasks, UniVS takes video frames and visual/text prompts as input, and the inference process is illustrated in the yellow boxes of Fig. 4. UniVS can process multiple targets simultaneously. First, the image encoder transforms the first frame into multi-scale image embeddings. Consequently, the prompt encoder converts visual/text prompts of the target into prompt tokens. In our design, each target has its dedicated memory pool to store associated prompt tokens, and its prompt query is obtained by averaging the tokens in the memory pool. These queries are used by the mask decoder to predict the masks of targets in the current frame, which are then used as visual prompts of targets and fed back to the prompt encoder, thereby updating the target’s memory pool with new prompt information. In short, UniVS utilizes the prompt information of target objects stored in the memory pool to identify and segment the targets in subsequent frames, eliminating the cumbersome post-matching step in other unified models like SEEM [seem] and UNINEXT [uninext], where targets need to be filtered out from all predicted entities.
@@
For category-specified VS tasks, UniVS adopts a periodic object detection strategy and transforms the segmentation into a prompt-guided target segmentation problem. The detailed process is depicted in the light green box in Fig. 4. First, UniVS employs learnable queries to identify all entity masks presented in the first frame, then employs non-maximum suppression (NMS) and classification thresholding to filter out redundant masks and those with low classification confidence. The remaining target objects also serve as their visual prompts, with which UniVS employs the prompt-guided target segmentation stream to directly predict their masks in the following frames, eliminating the need of cross-frame entity matching in previous methods. Furthermore, to identify newly appearing objects in subsequent frames, UniVS performs target detection for every a few frames using a learnable query and compares them to previously detected objects stored in a memory pool. We use the bi-Softmax approach [pang2021quasi] to distinguish between old and new objects in the video.

### v2

@@ 1 Introduction
Video segmentation (VS) partitions a video sequence into different regions or segments, facilitating many applications such as semantic-guided video restoration [39, 81, 71], video generation [59, 66], video editing and augmented reality [75, 97, 20], etc. VS tasks can be divided into two groups: category-specified VS and prompt-specified VS. The former focuses on segmenting and tracking entities from a predefined set of categories. Typical tasks along this line include video instance [87, 63], semantic [57] and panoptic segmentation [29, 56] (VIS/VSS/VPS), where the object category information needs to be specified. Another group focuses on identifying and segmenting specific targets throughout the video, where visual prompts or textual descriptions of the targets need to be provided. Prompt-specified VS tasks include video object segmentation (VOS) [61, 84, 16, 2], Panoptic VOS (PVOS) [85] and referring VOS (RefVOS) [68]. Each of above VS tasks has established its own protocol for dataset annotation, as well as model evaluation.
VS has a close relationship with image segmentation (IS) [18, 48, 100, 31, 13, 83, 54, 28, 101, 60], which has similar types of segmentation tasks to VS. In the past decades, the model performance on each individual IS/VS task has been significantly improved, and many well-known network architectures have been developed [22, 3, 5, 52, 67, 88, 7, 54, 91, 93, 37, 72, 12, 80, 26, 19]. Some architectures [96, 9, 33, 103, 44, 69, 43, 95] have also been proposed to deal with multiple segmentation tasks; however, these architectures require separate training and inference for different tasks because of the different annotations and evaluation protocols. Fortunately, the recent advancements in vision-language models [35, 65, 53, 36, 99, 27, 98, 49, 46, 50, 90] align and harmonize the multimodal feature representations, which bridge the labeling gaps across various IS/VS tasks. As a result, some unified segmentation models [64, 104, 34] have emerged to process multiple segmentation tasks simultaneously, which can be jointly trained on different datasets and tasks. These methods generally convert prompt-guided segmentation into the category-specified segmentation problem [22, 102, 9], which first predicts masks for all potential entities per frame and then utilizes a post-processing matching to find the target. For example, SEEM [104] captures prompt information of the target via concatenating learnable queries and prompt features as keys and values in the self-attention layer of the mask decoder, showing good versatility in various IS tasks.
@@
Compared to IS, the additional challenge of VS lies in the temporal consistency [73] of segmentation across frames in a video sequence. Existing unified models [86, 1] for VS tasks are mostly inspired by the unified IS models. They segment the video sequence frame by frame, and then use a similarity matching step to associate common objects or find the targets for category-specified and prompt-specified VS tasks, respectively. For example, UNINEXT [86] is specially designed for object-centric IS and VS tasks. Though UNINEXT performs very well in several aspects, it becomes ineffective when segmenting ‘stuff’ entities, such as ‘sky’. To adapt to different VS tasks, TarVIS [1] subdivides the learnable queries within the mask decoder into four groups: semantic, class-agnostic instance, object and background queries. However, TarVIS falls short in encoding linguistic information, hindering it from resolving language-guided VS tasks.
@@
To alleviate the above issues, we propose a novel unified VS architecture, namely UniVS, by using prompts as queries. For each target of interest, UniVS averages the prompt features from previous frames as its initial query. A target-wise prompt cross-attention (ProCA) layer is introduced in the mask decoder to integrate comprehensive prompt features stored in the memory pool. The initial query and the ProCA layer play a crucial role in the explicit and accurate decoding of masks. On the other hand, by taking the predicted masks of entities from previous frames as their visual prompts, UniVS can convert different VS tasks into the task of prompt-guided target segmentation task, eliminating the heuristic inter-frame matching. The overall process of UniVS is depicted in Fig. 2(b). UniVS not only unifies the different VS tasks (see Fig. 1) but also naturally achieves universal training and testing, resulting in robust performance across different scenarios. It shows a commendable balance between performance and universality on 10 challenging VS benchmarks, covering VIS, VSS, VPS, VOS, RefVOS and PVOS tasks. To the best of our knowledge, UniVS is the first work which can unify all the existing VS tasks successfully in a single model.
@@ 2 Related Work
Category-specified VS tasks includes video instance [87], semantic [57] and panoptic segmentation [56] tasks, which aim to correctly partition the regions and label them with specific categories or open-world vocabulary[92, 41, 19]. Building on representative query-based image detection and segmentation methods [4, 102, 40, 9], recent approaches [37, 72, 77, 80, 25, 26, 38, 24, 43, 30, 44] focus on efficiently encoding both short-term and medium-term temporal information. Some methods [77, 80, 26, 38, 45] utilize the consistency of the relative relationship between objects in a short period of time to associate entities across frames, while the latest state-of-the-arts [25, 24, 44, 43, 95] design learnable trackers between multiple video clips to better learn entity motion information over short to medium time periods, thus maintaining more temporally consistent segmentation results. This continual learning strategy also bridges the gap between training and inference.
Prompt-specified VS tasks consist of video object segmentation (VOS) [84, 62, 16, 2], panoptic VOS [85] and referring VOS [68], which aim to re-identify and segment the target objects with visual/text prompts in a video. Recent offline models [12, 89, 78, 93, 42, 55, 10] focus more on designing long-term information propagation modules to transfer previous image features and corresponding masks to the target frame to predict masks. This helps more precisely identify and track the movement trajectory of objects throughout the entire video sequence (more than 200 frames), but also makes the network architecture heavy [89, 79]. Due to the slow speed of offline models, online models [76, 94, 79] come back into focus, aiming to keep the right balance between speed and performance.
Unified VS models [6, 44, 69, 43, 95, 86, 1, 11, 79] have emerged recently. TubeLink [43] and DVIS [95] employ a single framework to accomplish different category-specified VS tasks, but they still need to be trained separately for each task and cannot be used for prompt-specified VS tasks. UniRef [79] uses a model to unify prompt-specified VS tasks, but it cannot handle category-specified VS tasks. To handle more types of VS tasks, TarVIS [1] decouples learnable queries of mask decoder into four groups. But, it fails to solve RefVOS due to the lack of language encoding capability. UNINEXT [86] is a unified object-centric segmentation model on images and videos, demonstrating universal applicability in several aspects. Unfortunately, it cannot handle entities with ‘stuff’ categories.
@@ 3 Methodology
UniVS contains of three main modules: the Image Encode, the Prompt Encoder and the Unified Video Mask Decoder, as depicted in Fig. 3. The Image Encoder transforms RGB images to the feature tokens, while the Prompt Encoder translates raw visual/text prompts into the prompt embeddings. The Unified Video Mask Decoder explicitly decodes masks for any entity or prompt-guided target in the video.
@@ 3.1 Image and Prompt Encoders
Image encoder contains a backbone and a pixel decoder [9]. The backbone maps the RGB image $X\in R^{3\times H\times W}$ into multi-scale features, and the pixel decoder further fuses features across scales to enhance the representation: $\{F_{s}\}_{s=1}^{4}=\text{ImEnc}(X),$ where $F_{s}\!\in\!R^{C\times H_{s}\times W_{s}}$ is the $s$-th scale image embeddings, $H_{s},W_{s}$ are its height and width and $C$ is the number of channels. The resolutions of multi-scale feature maps are 1/32, 1/16, 1/8 and 1/4 of that of the input image, respectively.
Prompt encoder converts the input visual/text prompts into the prompt embeddings. The visual prompts can be clicked points, boxes, masks and scribbles, etc. For the $i$-th target, we use $X_{i}^{*}\in R^{l_{i}\times 3}$ and $y_{i}^{*}\in R^{l_{i}}$ to represent its prompt-specified pixels and corresponding segment IDs, where $l_{i}$ is the total number of prompt-specified pixels. To convert the visual prompts into image embeddings, we adopt the Visual Sampler strategy proposed in SEEM [104]. It samples $l^{*}$ points from the prompt-specified pixels for each target, and extracts point features from the 3rd scale image embeddings $F_{3}$ as its visual prompt embeddings:
@@
The language prompt can be category names, such as ‘person’, and textual expressions, such as ‘a person is skateboarding’. Following [14, 65], we feed a category name or an expression into a tokenizer to get its string tokens $L_{i}^{*}\in R^{l^{*}\times C^{t}}$, which are input into the CLIP text encoder to obtain the text embeddings. We then introduce a single cross-attention layer to achieve language-image embeddings interaction, where the query is text embeddings, the keys and values are flattened multi-scale image embeddings $F\in R^{C\times(\sum_{s=1}^{3}H_{s}W_{s})}$. This process can be formulated as:
@@ 3.2 Unified Video Mask Decoder
where $P_{i}^{*}$ and $y_{i}^{*}$ are prompt features and associated segment IDs for the $i$-th target, and $M_{i}^{t}$ is its predicted masks in the $t$-th frame. $N_{p}$ and $V$ are the total numbers of provided targets and frames in the video, respectively. To achieve our goal, we adapt the mask decoder of Mask2Former [9, 6], which was initially designed for generic segmentation tasks with a set of learnable queries, by introducing a side stream that takes the mean of prompt features as input queries. As illustrated in the right yellow area of Fig. 3, our proposed unified video mask decoder comprises four key components: target-wise prompt cross-attention layer, an image cross-attention layer, a separated self-attention layer and a feed-forward network (FFN, which is omitted in Fig. 3).
@@ 4 Training and Inference
4.1 Training Stages
Training Losses consist of three terms: pixel-wise mask supervision loss, classification loss, and ReID loss. The training process of UniVS consists of three stages: image-level training, video-level training and long video fine-tuning. In the first stage, UniVS is trained on multiple image segmentation datasets, pretraining the model with image-level annotations for a good visual representation. In the second stage, we feed a short clip of three frames to the pretrained model, and fine-tune it on video segmentation datasets to perceive entity changes over a short period of time. In the third stage, we employ long video sequences of more than five frames to further fine-tune the unified video mask decoder, encouraging it to learn more discriminative features and trajectory information over a longer time period. To optimize memory usage, we freeze the backbone weights in the last two stages and further freeze the pixel decoder in the final stage. In each iteration, all samples within a batch come from the same dataset. We found that this sampling strategy can make the training more stable compared with the mixed sampling from different datasets.
@@ 4.2 Unified Streaming Inference Process
For prompt-specified VS tasks, UniVS takes video frames and visual/text prompts as input, and the inference process is illustrated in the yellow boxes of Fig. 4. UniVS can process multiple targets simultaneously. First, the image encoder transforms the first frame into multi-scale image embeddings. Consequently, the prompt encoder converts visual/text prompts of the target into prompt tokens. In our design, each target has its dedicated memory pool to store associated prompt tokens, and its prompt query is obtained by averaging the tokens in the memory pool. These queries are used by the mask decoder to predict the masks of targets in the current frame, which are then used as visual prompts of targets and fed back to the prompt encoder, thereby updating the target’s memory pool with new prompt information. In short, UniVS utilizes the prompt information of target objects stored in the memory pool to identify and segment the targets in subsequent frames, eliminating the cumbersome post-matching step in other unified models like SEEM [104] and UNINEXT [86], where targets need to be filtered out from all predicted entities.
@@
For category-specified VS tasks, UniVS adopts a periodic object detection strategy and transforms the segmentation into a prompt-guided target segmentation problem. The detailed process is depicted in the light green box in Fig. 4. First, UniVS employs learnable queries to identify all entity masks presented in the first frame, then employs non-maximum suppression (NMS) and classification thresholding to filter out redundant masks and those with low classification confidence. The remaining target objects also serve as their visual prompts, with which UniVS employs the prompt-guided target segmentation stream to directly predict their masks in the following frames, eliminating the need of cross-frame entity matching in previous methods. Furthermore, to identify newly appearing objects in subsequent frames, UniVS performs target detection for every a few frames using a learnable query and compares them to previously detected objects stored in a memory pool. We use the bi-Softmax approach [58] to distinguish between old and new objects in the video.